# Fourier Decay of Absolutely and Hölder Continuous Functions with Infinitely or Finitely Many Oscillations

### v3

@@ 2 Definitions
Thus the factorial polynomials behave similarly with respect to forward differences as typical polynomials do with respect to differentiation. The three asymptotic equivalences mentioned earlier are utilized to prove the following Lemma.
@@ 3 Known results on smoothness and Fourier decay
where $0<\mu<1$. It is stated in [7, p. 42] that the Fourier coefficients $c_{k}(g_{\mu})=O(1/|k|^{1+\mu})$, although one can criticize that partial integration was used twice in the proof, since the second application gives a divergent integral. For this reason I will go through the correct proof. The integral is divided into two parts after the partial integration and change of variables $x=t/k$, $\mathrm{d}x=\mathrm{d}t/k$ (here we consider $k>0$)
@@
Similar decay rate is of course achieved for some subsets of $BV[0,T]$, which contains for example piecewise continuous or monotone functions. These kind of special function spaces are also utilized to estimate the errors when estimating Fourier transforms and coefficients with the DFT for example in [12], [5, pp. 184–211].
@@
The previous result can be easily utilized to obtain an interesting result: suppose that the Fourier coefficients of a $T$-periodic function $f$ decay like $c_{k}(f)=O(1/|k|^{1+m+\mu})$, with $\mu\in\left(0.5,1\right)$, then $f\in W_{m+1}^{1}[0,T]$ as well. This follows from the following estimate and Lemma 3.11
@@ 4 Proof of the main result
Perhaps a little simpler proof would utilize that fact that we could also replace the infinite sums with integrals, since they are asymptotically equal in these cases. A similar result for multidimensional Fourier series can be found in [22, p. 178], although in one dimension it only states that the decay rate $c_{k}(f)=O(1/|k|^{1+m+\mu})$ implies $C_{\alpha}[0,T]$ for all $\alpha<\mu$.
@@ 5 Chirps
In [1, p. 331] it is stated that $f_{\alpha,\beta}$ is of bounded variation iff $\alpha>\beta$ and we will utilize this result now. Thus, let us suppose that $\alpha>\beta$. Due to the Lebesgue decomposition of a function of bounded variation in Lemma 2.6, we can deduce then that $\varphi$ and $\chi$ in this decomposition are both zero (since $f_{\alpha,\beta}$ is infinitely differentiable a.e. on $[0,L]$ and it is continuous on $[0,L]$) and thus $f_{\alpha,\beta}\in AC[0,L]$ iff $\alpha>\beta$.
@@ 6 Application to error analysis of numerical Weyl fractional derivatives
In this section we take a quick look at how Theorem 1.1 relates to fractional calculus and its error analysis as derived by the author in [28]. The following fractional differintegral operator is a generalization of the one studied by H. Weyl in [29].
@@
although the term $F_{N/2}$ is neqlected in the analysis. A bit more generalized result featuring the same decay rate is also given in eqs. (14) and (15) in [32]. An example of the result (70) for $\sigma=1$ was also given in [31], in which case for a function $f\in H^{s}(0,T)$, $s>1$, the $L^{2}$ error of the numerical first derivative with the Fourier method is
@@ 7 Extension to Fourier transforms
The theory and application of Fourier transforms is simplified considerably if we extend its definition to tempered distributions $S^{\prime}(\mathbb{R})$. These are linear continuous functionals from the space of Schwartz functions $S(\mathbb{R})$, which are also called rapidly decaying functions, to $\mathbb{C}$. As an example of the types of generalized functions in $S^{\prime}(\mathbb{R})$, we remind that given any polynomial $q$ and a function $g\in L^{p}(\mathbb{R})$ for any $1\leq p\leq\infty$, their product $pg\in S^{\prime}(\mathbb{R})$.
A generalized function $f\in S^{\prime}(\mathbb{R})$ has generalized derivatives of all order and they all have Fourier transforms as tempered distributions.
@@
Note that necessarily $C=\lim_{t\rightarrow-\infty}f(t)$ in (82). This definition is almost equal to Sobolev space $W_{1}^{1}(\mathbb{R})$, which means that the function and its first generalized derivative are in $L_{1}(\mathbb{R})$ and that (82) applies. Note that the two requirements $f,f^{\prime}\in L_{1}(\mathbb{R})$ of the space $W_{1}^{1}(\mathbb{R})$ guarantee that $\lim_{|t|\rightarrow\infty}f(t)=0$. We will not need the integrability of $f$ in the Theorem that follows and thus we first state it without Sobolev spaces.
@@
Proved using generalized derivatives in [24, pp. 33 – 34]. ∎
@@
This was proved for functions with polynomial decay in [39] and [40], so we generalize the proof here.
@@
For $\mu\in(0,1]$, the function $e^{-\pi t^{2}}|t|^{\mu}$ satisfies the requirements of Corollary 7.14. Thus, we know that $\mathcal{F}\left\{e^{-\pi t^{2}}|t|^{\mu}\right\}(\nu)=O(1/|\nu|^{1+\mu})$ as $|\nu|\rightarrow\infty$. The convolution Theorem for rapidly decaying Schwartz functions $S(\mathbb{R})$ and tempered distributions [8, pp. 170 – 171] can be used to decompose this (since $e^{-\pi t^{2}}\in S(\mathbb{R})$ and $|t|^{\mu}\in S^{\prime}(\mathbb{R})$ )
@@
$=\left(e^{-\pi\nu^{2}}\ast\frac{-2\sin\left(\frac{\pi\mu}{2}\right)\Gamma(\mu+1)}{|2\pi\nu|^{1+\mu}}\right)(\nu),$ (98)
where we substituted the Fourier transforms of the two functions. Because these transforms are known, we could in this case also apply Lemma 7.17 to prove that the decay rate of $\mathcal{F}\left\{e^{-\pi t^{2}}|t|^{\mu}\right\}(\nu)$ is governed by the decay rate of $\mathcal{F}\left\{|t|^{\mu}\right\}(\nu)$. The formula for $\mathcal{F}\left\{|t|^{\mu}\right\}$ and other similar ones is proved, for example, in [38, pp. 170 – 173]. Since the distributional Fourier transform $\mathcal{F}\left\{|t|^{\mu}\right\}$ also decays like $O(1/|\nu|^{1+\mu})$, it is likely that a generalization of Theorem 7.12 to some tempered distributions with enough smoothness is possible.
@@ Acknowledgments
I wish to thank Jukka Kemppainen for his comments on an early version of this manuscript. I also thank professor Valery Serov for his excellent courses on the theory of Fourier series and the Fourier transform organized at the University of Oulu.
All web addresses retrieved on 18.09.2025.

### v4

@@ 2 Definitions
Thus the factorial polynomials behave similarly with respect to forward differences as typical polynomials do with respect to differentiation. The three asymptotic equivalences mentioned earlier are utilised to prove the following Lemma.
@@ 3 Known results on smoothness and Fourier decay
where $0<\mu<1$. It is stated in [7, p. 42] that the Fourier coefficients $c_{k}(g_{\mu})=O(1/|k|^{1+\mu})$, although one can criticise that partial integration was used twice in the proof, since the second application gives a divergent integral. For this reason I will go through the correct proof. The integral is divided into two parts after the partial integration and change of variables $x=t/k$, $\mathrm{d}x=\mathrm{d}t/k$ (here we consider $k>0$)
@@
Similar decay rate is of course achieved for some subsets of $BV[0,T]$, which contains for example piecewise continuous or monotone functions. These kind of special function spaces are also utilised to estimate the errors when estimating Fourier transforms and coefficients with the DFT for example in [12], [5, pp. 184–211].
@@
The previous result can be easily utilised to obtain an interesting result: suppose that the Fourier coefficients of a $T$-periodic function $f$ decay like $c_{k}(f)=O(1/|k|^{1+m+\mu})$, with $\mu\in\left(0.5,1\right)$, then $f\in W_{m+1}^{1}[0,T]$ as well. This follows from the following estimate and Lemma 3.11
@@ 4 Proof of the main result
Perhaps a little simpler proof would utilise that fact that we could also replace the infinite sums with integrals, since they are asymptotically equal in these cases. A similar result for multidimensional Fourier series can be found in [22, p. 178], although in one dimension it only states that the decay rate $c_{k}(f)=O(1/|k|^{1+m+\mu})$ implies $C_{\alpha}[0,T]$ for all $\alpha<\mu$.
@@ 5 Chirps
In [1, p. 331] it is stated that $f_{\alpha,\beta}$ is of bounded variation iff $\alpha>\beta$ and we will utilise this result now. Thus, let us suppose that $\alpha>\beta$. Due to the Lebesgue decomposition of a function of bounded variation in Lemma 2.6, we can deduce then that $\varphi$ and $\chi$ in this decomposition are both zero (since $f_{\alpha,\beta}$ is infinitely differentiable a.e. on $[0,L]$ and it is continuous on $[0,L]$) and thus $f_{\alpha,\beta}\in AC[0,L]$ iff $\alpha>\beta$.
@@ 6 Application to error analysis of numerical Weyl fractional derivatives
In this section we take a quick look at how Theorem 1.1 relates to fractional calculus and its error analysis as derived by the author in [28]. The following fractional differintegral operator is a generalisation of the one studied by H. Weyl in [29].
@@
although the term $F_{N/2}$ is neqlected in the analysis. A bit more generalised result featuring the same decay rate is also given in eqs. (14) and (15) in [32]. An example of the result (70) for $\sigma=1$ was also given in [31], in which case for a function $f\in H^{s}(0,T)$, $s>1$, the $L^{2}$ error of the numerical first derivative with the Fourier method is
@@ 7 Extension to Fourier transforms
The theory and application of Fourier transforms is simplified considerably if we extend its definition to tempered distributions $S^{\prime}(\mathbb{R})$. These are linear continuous functionals from the space of Schwartz functions $S(\mathbb{R})$, which are also called rapidly decaying functions, to $\mathbb{C}$. As an example of the types of generalised functions in $S^{\prime}(\mathbb{R})$, we remind that given any polynomial $q$ and a function $g\in L^{p}(\mathbb{R})$ for any $1\leq p\leq\infty$, their product $pg\in S^{\prime}(\mathbb{R})$.
A generalised function $f\in S^{\prime}(\mathbb{R})$ has generalised derivatives of all order and they all have Fourier transforms as tempered distributions.
@@
Note that necessarily $C=\lim_{t\rightarrow-\infty}f(t)$ in (82). This definition is almost equal to Sobolev space $W_{1}^{1}(\mathbb{R})$, which means that the function and its first generalised derivative are in $L_{1}(\mathbb{R})$ and that (82) applies. Note that the two requirements $f,f^{\prime}\in L_{1}(\mathbb{R})$ of the space $W_{1}^{1}(\mathbb{R})$ guarantee that $\lim_{|t|\rightarrow\infty}f(t)=0$. We will not need the integrability of $f$ in the Theorem that follows and thus we first state it without Sobolev spaces.
@@
Proved using generalised derivatives in [24, pp. 33 – 34]. ∎
@@
This was proved for functions with polynomial decay in [39] and [40], so we generalise the proof here.
@@
For $\mu\in(0,1]$, the function $e^{-\pi t^{2}}|t|^{\mu}$ satisfies the requirements of Corollary 7.14. Thus, we know that $\mathcal{F}\left\{e^{-\pi t^{2}}|t|^{\mu}\right\}(\nu)=O(1/|\nu|^{1+\mu})$ as $|\nu|\rightarrow\infty$. The convolution Theorem for rapidly decaying Schwartz functions $S(\mathbb{R})$ and tempered distributions [8, pp. 170 – 171] can be used to decompose this (since $e^{-\pi t^{2}}\in S(\mathbb{R})$ and $|t|^{\mu}\in S^{\prime}(\mathbb{R})$ ) when $\mu\in(0,1)$
@@
$=e^{-\pi\nu^{2}}\ast\frac{-2\sin\left(\frac{\pi\mu}{2}\right)\Gamma(\mu+1)}{|2\pi\nu|^{1+\mu}},$ (98)
where the generalised function $\mathcal{F}\left\{|t|^{\mu}\right\}(\nu)$ is defined via its Hadamard finite part integral. Therefore, even though the Fourier transforms are known, the latter is not in $L_{1}(\mathbb{R})$ and Lemma 7.17 could not be directly used to prove that the decay rate of $\mathcal{F}\left\{e^{-\pi t^{2}}|t|^{\mu}\right\}(\nu)$ is governed by the decay rate of $\mathcal{F}\left\{|t|^{\mu}\right\}(\nu)$. The formula for $\mathcal{F}\left\{|t|^{\mu}\right\}$ and other similar ones is proved, for example, in [38, pp. 170 – 173] or in Mark Viola’s answer from 2021 in [41]. Since the distributional Fourier transform $\mathcal{F}\left\{|t|^{\mu}\right\}$ also decays like $O(1/|\nu|^{1+\mu})$, it is likely that a generalisation of Theorem 7.12 to some tempered distributions with enough smoothness is possible.
@@ Acknowledgments
I wish to thank Jukka Kemppainen for his comments on an early version of this manuscript. I also thank professor Valery Serov for his excellent courses on the theory of Fourier series and the Fourier transform organised at the University of Oulu.
All web addresses retrieved on 18.03.2026.